OpenStudy (anonymous):

ABCD is a rectangle. Find the length of each diagonal. ac= 3y/5 bd= 3y-4

OpenStudy (anonymous):

if ABCD is a rectangle then the diagonals are equal. Are the expressions u posted thw length and width?

OpenStudy (anonymous):

no just length of the lines

OpenStudy (anonymous):

they cant be diagonals of a rectangle because all rectangles have equal diagonals. maybe it is a parallelogram?

OpenStudy (anonymous):

yes it is its just that abcd is the rectangle just solve for y

OpenStudy (anonymous):

\[3y/5=3y-4\] \[3y=15y-20\] \[12y=20\] \[y=5/3\] \[ac=1=bd=1\]